OpenStudy (anonymous):

The price of a suit decreased from $280 to $182 during a clearance sale. What is the percent of decrease in the price? Round to the nearest percent, if necessary. A. 98% B. 65% C. 54% D. 35%

OpenStudy (anonymous):

The price decreased by $98. The original price was $280 so it is the percentage decrease of this figure we are looking for. The calculation is:- 98/280 x 100 = 9800/280 = 35 so the decrease is 35%

OpenStudy (anonymous):

You know they have this in google lol http://answers.yahoo.com/question/index?qid=20110202113925AAZ2Urf